Schedule of Cash Collections of Accounts Receivable Pet Supplies Inc., a pet wholesale supplier, was organized on January 1. Projected sales for each of the first three months of operations are as follows:

January $300,000

February 500,000

March 750,000

All sales are on account. Seventy-five percent of sales are expected to be collected in the month of the sale, 20% in the month following the sale, and the remainder in the second month following the sale.

Prepare a schedule indicating cash collections from sales for January, February, and March. Enter all amounts as positive numbers.

Pet Supplies Inc. Schedule of Cash Collections from Sales For the Three Months Ending March 31
January February March
January Sales on Account
Collected in January $
Collected in February $
Collected in March $
February Sales on Account
Collected in February $
Collected in March $
March Sales on Account
Collected in March $
Total Cash Collected $ $ $
Answer #1

Ans.

Given:

statement showing sales on account

Month Amount($)
January $300,000
February $500,000
March $750,000

Pet Supplies Inc. Schedule of Cash Collections from Sales For the Three Months Ending

Particulars January February March
January Sales on Account
Collected in January $225,000
Collected in February $60,000
Collected in March $15,000
February Sales on Account
Collected in February $375,000
Collected in March $100,000
March Sales on Account
Collected in March $562,500
Total Cash Collected $225,000 $435,000 $677,500

Working :

Statement showing collection of cash for the month of January

Particulars collection (%) Amount($)
Collection in January 75% = $300,000*75%= $225,000
Collection in February 20% = $300,000*20%=$60,000
Collection in March 5% (Residual) = $300,000*5%= $15,000

Statement showing collection of cash for the month of February

Particulars Collection (%) Amount ($)
Collection in February 75% = $500,000*75%= $375,000
Collection in March 20% = $500,000*20%= $100,000

Statement showing collection of cash for the month of March

Particulars Collection (%) Amount ($)
Collection in March 75% = $750,000*75% =$562,500